OpenStudy (anonymous):

Rewrite with only sin x and cos x. cos 3x

OpenStudy (anonymous):

cos x - 4 cos x sin^2x -sin^3x + 2 sin x cos x -sin^2x + 2 sin x cos x 2 sin^2x cos x - 2 sin x cos x

OpenStudy (dumbcow):

\[\cos(3x) = \cos(2x+x)\] \[=\cos(2x)\cos x - \sin(2x)\sin x\] \[=(1-2\sin^{2}x)\cos x - 2\sin^{2} x \cos x\] \[\cos x - 4\sin^{2}x \cos x\]

OpenStudy (anonymous):

Thats what i got too but its not one of the answer choices so im confused...?

OpenStudy (dumbcow):

look at option 1 :)

OpenStudy (anonymous):

oh its just in a different order... oops :p thanks @dumbcow